OpenStudy (anonymous):

what is the slope of the line that contains the points (-7,5) (-2,3)? choices: -5/2, -2/5, 2/5, 5/2

OpenStudy (anonymous):

slope formula: \[\frac{ y2-y1 }{ x2-x1 }\] (x1,y1)(y1,y2)

OpenStudy (anonymous):



OpenStudy (anonymous):

(-7,5)(-2,3) and then fill it to the formula \[\frac{ \left( -2-(-7) \right) }{ \left( 3-5 \right) }\]

OpenStudy (anonymous):

find the answer.. :D

OpenStudy (anonymous):

just give me the answer

OpenStudy (anonymous):

thanks dalia lam dude